OpenStudy (anonymous):

The average of two real numbers is 12.75 and their product is 162.5. What are the numbers? a.12.5, 13 b. 12, 13 c. 10.8,15 d. 9,18.1

OpenStudy (johnweldon1993):

\[\large \frac{a + b}{2} = 12.75\] \[\large a \times b = 162.5\] Solve the system of equations